I have a word problem that I'm thinking does not have enough info. to solve. The problem is:
"Find the length of the sides of a square if the number of square meters in its area is 1.5 times the number of meters in its perimeter."

Answer by ankor@dixie-net.com(22740)

"Find the length of the sides of a square if the number of square meters in its area is 1.5 times the number of meters in its perimeter."
:
Let x = the side of the square
then
x^2 = it's area
and
4x = it's perimeter
:
Write an equation for the given statement.
"the number of square meters in its area is 1.5 times the number of meters in its perimeter."
x^2 = 1.5(4x)
x^2 = 6x
x^2 - 6x = 0
Factor
x(x-6) = 0
two solutions
x = 0
and
x = 6 meters is the answer we want
:
:
See if that works out
6^2 = 1.5(4*6)